OpenStudy (anonymous):

x^2+y^2=169 is y a function of x and how do you know?

OpenStudy (imstuck):

If y is a function of x, for every x input there is exactly one y output. That is the definition of a function. Here, you could put in both a positive and negative of a number because both positive and negative numbers squared yield the same number. This is not a function. It doesn't pass the vertical line test, either. That's the first clue.

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ x^2+y^2=169}\) \(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ (x-0)^2+(y-0)^2=13^2}\) \(\LARGE\color{white}{ \rm │ }\)

OpenStudy (solomonzelman):

a circle is not a function.

OpenStudy (solomonzelman):

IMStuck is right, it fails to pass vertical test...